# Fast Adaptive Anti-Jamming Channel Access via Deep Q Learning and Coarse-Grained Spectrum Prediction

### v2

@@ 5.2 Results
We first investigate the performance of the proposed fast adaptive anti-jamming approach under traditional jamming attacks. Specifically, we assume that the intelligent jammer is dormant, while the fixed-mode jammer could launch traditional jamming attacks with a fixed jamming pattern, such as frequency sweeping jamming or comb jamming, etc. In this case, the anti-jamming problem could be simplified to a Markov decision problem, and the proposed approach is compared with the traditional pseudo-random sequence based FH approach Navda et al. (2007) and the emerging DRL-based FH approach Wang et al. (2020), with the results shown in Fig. 7. It is noted that each curve is averaged over 5 independent trials. Under different jamming modes, the throughput of the pseudo-random sequence-based approach is maintained at approximately 65%, while the throughput of the DRL-based approach converges to 97%. This is because, under a fixed jamming mode, the RL-based anti-jamming approach could effectively learn the jammer’s behavior and generate corresponding anti-jamming strategies. Additionally, the proposed approach could achieve a throughput close to 100% and demonstrates a substantially faster convergence rate in model training compared to the DRL-based approach, requiring approximately 80% fewer training episodes. This improvement is attributed to the additional information gain that the coarse-grained spectrum prediction provides to the DRL.
Then, we compare the anti-jamming performance of the proposed fast adaptive anti-jamming approach with several existing approaches under the combined influence of a fixed-mode jammer and a DRL-based jammer. The fixed-mode jammer could launch sweeping jamming signals Pirayesh and Zeng (2022) with a sweeping speed of 500MHz/s, effectively covering the entire communication band through linear frequency changes. In the experiment, the following four comparison approaches are considered:
  1. Traditional frequency hopping (FH) Navda et al. (2007): The working frequency of the legitimate user is determined by a predefined FH pattern with finite length.
  2. DRL-based approach Wang et al. (2020): The jammer is regarded as a component of the anti-jamming environment, and the DQN based approach could be introduced for combating jamming attacks.
  4. Opponent modeling based approach Li et al. (2023a): This method adopts the minimax DQN to approximate the user’s utility while the imitation learning method is employed to infer the intelligent jammer’s strategy. The OMAIJ algorithm could adapt to the dynamic behavior of the intelligent jammer and identify the BR instead of aiming for the NE.
@@
(a) Comparison of different anti-jamming approaches.
(b) Comparison of the proposed approach under different jamming update steps.
(c) Ablation experiment.
Figure 8: Anti-jamming performance of the proposed approach when facing the DRL-based intelligent jamming.
@@
The normalized throughputs of these anti-jamming approaches over 300 training episodes are shown in Fig. 8LABEL:sub@8a, where each curve is averaged over 5 independent trials. It is evident that our approach outperforms the traditional FH approach, and could achieve about 10% improvements over the NE-based approach and the DRL-based approach. The traditional frequency hopping approach makes it difficult for the RL-based jammer to devise an effective response, but it does not enhance the legitimate user’s anti-jamming performance. What’s more, the learning based anti-jamming algorithms could increase the user’s normalized throughput by learning the jammers’ dynamics, and we find that the anti-jamming performance of the NE-based and DRL-based approaches is comparable. The proposed approach could achieve comparable anti-jamming performance to the opponent modeling based approach. Meanwhile, the proposed approach also demonstrates a significant acceleration in the rate of convergence and could reduce the number of training episodes by up to 70%. Moreover, our approach adopts the coarse-grained spectrum prediction as the auxiliary task and does not need any knowledge of the jamming action or the jammer’s action space, which might not be accessible in practical anti-jamming scenarios.
@@
Furthermore, we investigate the anti-jamming performance of the proposed approach when facing the intelligent jammer with different update steps, and the results are shown in Fig. 8LABEL:sub@8b. The smaller update step implies that the jammer’s policy could be updated more frequently. As the jammer accelerates its updating process, the anti-jamming performance deteriorates. This is because both the coarse-grained spectrum prediction network and the Q-function estimation network find it more challenging to learn the rapidly changing jamming policy. But even if the intelligent jammer takes the same update frequency as the legitimate user (i.e., Step=1), the normalized throughput could still achieve 70%, which is also much higher than the NE-based approach. If the intelligent jammer adopts a fixed jamming policy (i.e., Step = infinity), the normalized throughput could converge to about 95%. Therefore, if the user could update its policy more rapidly, the proposed approach could perform better when facing learning-based jammers.
@@
Finally, we conducted an ablation experiment to confirm the effectiveness of the proposed approach. As shown in Fig. 8LABEL:sub@8c, the proposed approach is compared with the opponent modeling based approach (DPIQN) Hong et al. (2018) and the coarse-grained spectrum prediction based approach. The DPIQN algorithm also uses an auxiliary task for the opponent’s policy feature extraction and chooses the action with the maximum Q-value for each hop, while the coarse-grained spectrum prediction based approach chooses the channel with the minimum predicted discrete spectrum sample value for data transmission at the beginning of each hop. The results in Fig. 8LABEL:sub@8c show that the proposed approach exhibits the strongest anti-jamming performance and the fastest rate of convergence in model training, followed by the supervised learning based coarse-grained spectrum prediction approach, with the DPIQN-based Q-function estimation approach being the least effective. This demonstrates that the supervised learning based approach converges faster than the DPIQN algorithm, and the proposed approach could effectively leverage the features extracted from DRL as well as coarse-grained spectrum prediction for joint decision-making.
@@ 6 Conclusion
This paper proposes a novel fast adaptive anti-jamming channel access approach that leverages deep Q learning and coarse-grained spectrum prediction to enhance anti-jamming performance while reducing training episodes in wireless confrontations. We formulate the adversarial interactions as a Markov game and address the limitations of existing opponent modeling techniques that rely on the observation of opponent actions by employing a supervised learning based coarse-grained spectrum prediction. This prediction serves as an auxiliary task for learning the intelligent jammer’s strategy, and an updated training dataset employing the FIFO principle is adopted to ensure its responsiveness to dynamic environments. Numerical simulations indicate that the proposed approach demonstrates a reduction of up to 70% in training episodes compared to DRL-based approaches, along with a 10% improvement in throughput over Nash equilibrium strategies. This work contributes a robust, efficient solution for adaptive anti-jamming in complex, real-world scenarios.

### v3

@@ 5.2 Results
We first investigate the performance of the proposed fast adaptive anti-jamming approach under traditional jamming attacks. Specifically, we assume that the intelligent jammer is dormant, while the fixed-mode jammer could launch traditional jamming attacks with a fixed jamming pattern, such as sweeping jamming or comb jamming, etc. In this case, the anti-jamming problem could be simplified to a Markov decision process. The proposed approach is compared with the traditional pseudo-random sequence based FH approach Navda et al. (2007), the emerging DRL-based FH approach Wang et al. (2020), and the state-of-the-art (SOTA) DRL-based approach (labeled DRL based anti-jamming approach Li et al. (2023b)), with the results shown in Fig. 8. In our simulations, we employ a Rayleigh fading channel model to capture the time-varying characteristics of wireless channels in non-line-of-sight (NLOS) scenarios. The Rayleigh distribution could effectively characterize the statistical fluctuations of the received signal envelope in such environments, thereby providing a more accurate representation of realistic fading behavior and introducing appropriate variability in the received signal power over time. As mentioned in Li et al. (2020), the anti-jamming performance of the legitimate user could be evaluated using the normalized throughput, which is defined as the successful transmission rate in every $K$ time slots, i.e.,
$\mathcal{T}=\frac{1}{K}\sum_{k=1}^{K}\delta(\beta(k,f_{k}^{u})\geq\beta_{\text{th}}),$ (34)
where $\beta(k,f_{u}^{k})$ represents the SINR at the receiver during the $k$-th time slot, $\beta_{\text{th}}$ denotes the demodulation threshold, and $\delta(\cdot)$ is the indicator function. It is noted that each curve is averaged over 5 independent trials. Across different jamming modes, the throughput of the pseudo-random sequence-based approach remains around 65%, whereas both the DRL-based and labeled DRL-based approaches converge to approximately 97%. This occurs because, under fixed-mode jamming attacks, the RL-based anti-jamming approach could effectively learn the jammer’s behavior and generate corresponding anti-jamming strategies. Moreover, the proposed approach achieves a throughput close to 100% and demonstrates a substantially faster rate of convergence during model training compared to the SOTA DRL-based approach, reducing the required training episodes by approximately 80%. This improvement is attributed to the additional information gain that the coarse-grained spectrum prediction provides to the DRL.
@@
Then, we compare the proposed fast adaptive anti-jamming approach with several existing approaches under the combined influence of a fixed-mode jammer and a DRL-based jammer. The fixed-mode jammer could launch sweeping jamming signals Pirayesh and Zeng (2022) with a sweeping speed of 500MHz/s, effectively covering the entire communication band through linear frequency changes. In the experiment, the following five comparison approaches are considered:
  1. Traditional FH Navda et al. (2007): The working frequency of the user is determined by a predefined FH pattern with finite length.
  2. Standard DRL-based approach Wang et al. (2020): The jammer is regarded as a component of the anti-jamming environment, and the DQN based approach could be introduced for combating jamming attacks. Specifically, the network hyperparameters of the DRL-based anti-jamming approach are set according to Liu et al. (2018).
  3. Labeled DRL-based FH Li et al. (2023b): A SOTA DRL-based frequency hopping approach that employs soft labels rather than rewards to accelerate the rate of convergence during the training of the anti-jamming model.
  5. Opponent modeling based FH Li et al. (2023c): This method adopts the minimax DQN to approximate the user’s utility while the imitation learning model is employed to infer the intelligent jammer’s strategy. The OMAIJ algorithm could adapt to the dynamic behavior of the intelligent jammer and identify the BR instead of aiming for the NE.
@@
Figure 9: Comparison of different anti-jamming approaches under DRL-based intelligent jamming attacks.
@@
The normalized throughputs of these anti-jamming approaches over 300 training episodes are shown in Fig. 9, where each curve is averaged over 5 independent trials. It is evident that our approach outperforms the traditional FH approach, and could achieve about 10% improvements over the NE-based approach and the standard DRL-based approach. The traditional frequency hopping approach makes it difficult for the RL-based jammer to devise an effective response, but it does not enhance the legitimate user’s anti-jamming performance. What’s more, the learning based anti-jamming algorithms could increase the user’s normalized throughput by learning the jammers’ dynamics, and we find that the anti-jamming performance of the NE-based and standard DRL-based approaches is comparable. Compared to SOTA DRL-based frequency hopping approach Li et al. (2023b), our approach demonstrates a slight improvement in anti-jamming performance and reduces the number of training episodes by approximately 50%. The proposed approach could achieve comparable anti-jamming performance to the opponent modeling based approach. Meanwhile, the proposed approach also demonstrates a significant acceleration in the rate of convergence and could reduce the number of training episodes by up to 70%. Moreover, our approach adopts the coarse-grained spectrum prediction as the auxiliary task and does not need any knowledge of the jamming action or the jammer’s action space, which might not be accessible in practical anti-jamming scenarios.
Additionally, we use prediction error to quantify the accuracy of the coarse-grained spectrum prediction and further investigate the accuracy of this CNN-based prediction during the online training process. Specifically, the coarse-grained spectrum prediction error is defined as
$D(\hat{\bf{c}}_{n},{\bf{c}}_{n})=\frac{\sqrt{\frac{1}{M}\sum_{m=1}^{M}\left(\hat{c}_{n}^{m}-c_{n}^{m}\right)^{2}}}{\max\left\{{\bf{c}}_{n}\right\}-\min\left\{{\bf{c}}_{n}\right\}},$ (35)
where $\hat{c}_{n}^{m}$ denotes the $m$-th element of the predicted coarse-grained spectrum $\hat{\bf{c}}_{n}$, and $c_{n}^{m}$ denotes the $m$-th element of the ground-truth label ${\bf{c}}_{n}$. Fig. 10 shows the relationship between coarse-grained spectrum prediction error and anti-jamming performance during the online training process. It can be observed that the coarse-grained spectrum prediction error gradually decreases with the increasing number of training episodes, while the throughput of the legitimate user correspondingly improves. This is because the accurately-estimated coarse-grained spectra assist the joint decision function in identifying better anti-jamming strategies, thereby significantly improving the performance of the anti-jamming agent.
Figure 10: The relationship between coarse-grained spectrum prediction error and anti-jamming performance during the online training process.
@@
Furthermore, we investigate the anti-jamming performance of the proposed fast adaptive anti-jamming approach under a more realistic non-synchronous time-slotted framing structure through simulations, where an offset $o\in\left(0,T_{h}/2\right)$ is assumed due to the switching of anti-jamming actions and jamming actions. The results are shown in Fig. 11, where the offset is set to $o=\Delta t,3\Delta t,5\Delta t$, respectively. It can be seen that the misalignment of time slots has no significant impact on the performance of the proposed anti-jamming approach. Even with an offset of $o=5\Delta t$, the anti-jamming performance of the proposed approach still significantly outperforms the SOTA DRL-based anti-jamming approach Li et al. (2023b), which uses a synchronous time-slotted framing structure, thereby demonstrating the robustness of the proposed fast adaptive anti-jamming approach. Additionally, we also investigate the anti-jamming performance of the proposed approach when facing the intelligent jammer with different update steps, and the results are shown in Fig. 12. The smaller update step implies that the jammer’s policy could be updated more frequently. As the jammer accelerates its updating process, the anti-jamming performance deteriorates. This is because both the coarse-grained spectrum prediction network and the Q-function estimation network find it more challenging to learn the rapidly changing jamming policy. But even if the intelligent jammer takes the same update frequency as the legitimate user (i.e., Step=1), the normalized throughput could still achieve 70%, which is also much higher than the NE-based approach. If the intelligent jammer adopts a fixed jamming policy (i.e., Step = infinity), the normalized throughput could converge to about 95%. Therefore, if the user could update its policy more rapidly, the proposed approach could perform better when facing learning-based jammers.
Figure 11: The performance of the proposed fast-adaptive anti-jamming method with non-synchronous time-slotted framing structure.
Figure 12: Comparison of the proposed approach under different jamming update steps.
@@
Finally, we conducted an ablation experiment to validate the effectiveness of the proposed approach. As shown in Fig. 13, the proposed approach is compared with the opponent modeling based approach (DPIQN) Hong et al. (2018), the coarse-grained spectrum prediction based approach, and the general spectrum prediction based approach Pan et al. (2023). The DPIQN algorithm also utilizes an auxiliary task for the opponent’s policy feature extraction and selects the action with the maximum Q-value in each hop, while the spectrum prediction based approaches select the channel with the minimum predicted discrete spectrum sample value for data transmission at the beginning of each hop. The results in Fig. 13 demonstrate that the proposed approach exhibits the strongest anti-jamming performance and the fastest rate of convergence during model training, followed by the supervised learning based coarse-grained spectrum prediction approach and the general spectrum prediction based approach, with the DPIQN based Q-function estimation approach being the least effective. This indicates that the supervised learning based approach converges faster than the DPIQN algorithm, and the proposed approach could effectively leverage the features extracted from DRL, as well as coarse-grained spectrum prediction, for joint decision-making. It is noted that the proposed approach and the coarse-grained spectrum prediction based approach demonstrate a reduction of up to 70% in training episodes compared to the general spectrum prediction based approach. This is because, compared to the general spectrum prediction, the proposed coarse-grained spectrum prediction task is much simpler. This improvement stems from the simplicity of the proposed coarse-grained spectrum prediction task compared to general spectrum prediction.
Figure 13: Ablation experiment.
5.3 Computational Complexity Analysis
The computational complexity for the proposed fast adaptive anti-jamming approach is measured in floating-point operations (FLOPs). In the case of a CNN, the computational complexity of a Conv layer is defined as
$\text{FLOPs}_{\text{Conv}}=2C_{\text{in}}K^{2}\times(H\times W\times C_{\text{out}}),$ (36)
where $K$ is the kernel size, $H\times W$ is the size of the output feature map, $C_{\text{in}}$ and $C_{\text{out}}$ are the number of input channels and output channels, respectively.
The computational complexity of a FC layer is defined as
$\text{FLOPs}_{\text{FC}}=2I\times O,$ (37)
where $I$ and $O$ are the number of input and output neurons.
The required FLOPs for both the Q-function estimation network and the coarse-grained spectrum prediction network are $143.9\times 10^{6}$, and the complexity of the proposed Algorithm 1 is $287.8\times 10^{6}$. Graphics processing unit (GPU) could satisfy the computational demands of the proposed anti-jamming algorithm with joint DQN and coarse-grained spectrum prediction. In simulations, each training episode—including all channel sensing, access, and jamming operations—takes approximately 5 seconds when executed on an NVIDIA RTX 4090 GPU. The actual duration may vary depending on the specific hardware platform and the simulation configuration. Additionally, a comparison between the proposed fast adaptive anti-jamming channel access approach and several benchmarks in terms of computational complexity, rate of convergence, and anti-jamming performance is shown in Table 2. It can be observed that although the proposed approach incurs higher computational complexity than the standard DRL-based method Wang et al. (2020), labeled DRL-based method Li et al. (2023b), and NE-based method Zhu and Zhao (2022), due to the use of dual-branch CNNs and coarse-grained spectrum prediction, it achieves significantly higher throughput and a faster rate of convergence during training. Furthermore, compared to the opponent modeling-based method (i.e., OMAIJ Li et al. (2023c)) with similar computational complexity, our approach attains slightly better anti-jamming performance and converges more rapidly. Therefore, the tradeoff between complexity and performance is justified by the improved anti-jamming effectiveness and training efficiency.
Table 2: Comparison of computational complexity, rate of convergence, and anti-jamming performance
Method Computational Complexity (FLOPs) Rate of convergence (episodes) Throughput
[…]
@@ 6 Conclusion
This paper proposes a novel fast adaptive anti-jamming channel access approach that leverages deep Q learning and CG spectrum prediction to enhance anti-jamming performance while reducing training episodes in wireless confrontations. We formulate the adversarial interactions as an MG and address the limitations of existing opponent modeling techniques that rely on the observation of opponent actions by employing a supervised learning based CG spectrum prediction. This prediction serves as an auxiliary task for learning the intelligent jammer’s strategy, and an updated training dataset employing the FIFO principle is adopted to ensure its responsiveness to dynamic environments. Numerical simulations indicate that the proposed approach demonstrates a reduction of up to 70% in training episodes compared to DRL-based approaches, along with a 10% improvement in throughput over NE strategies. This work contributes a robust, efficient solution for adaptive anti-jamming in complex, real-world scenarios. In the future, it would be valuable to explore practical implementations of the proposed anti-jamming approach, which could perform joint DQN and CG spectrum prediction, using software-defined radio devices.